# Partial Functional Dynamic Backdoor Diffusion-based Causal Model

### v1

@@ 2 Preliminaries
Diffusion models (Sohl-Dickstein et al.,, 2015; Ho et al.,, 2020) approximate a target data distribution $q(x^{0})$ via a two-stage process. First, a fixed forward process gradually injects Gaussian noise into the data $x^{0}$ over $T$ steps. The distribution of the noisy data $x^{t}$ at step $t$ is a Gaussian: $q(x^{t}|x^{0})=\varphi(x^{t};\sqrt{\alpha_{t}}x^{0},(1-\alpha_{t})I)$, where ${\varphi}(x;\mu,\Sigma)$ denote the Gaussian density with mean $\mu$ and covariance $\Sigma$. Here, $\alpha_{t}:=\prod_{s=1}^{t}(1-\beta_{s})$ where $\beta_{s}$ is a predefined noise schedule at each time step $s$. As $t\to T$, $x^{T}$ converges to a standard Gaussian distribution.
Second, a learnable reverse process, parameterized by $\theta$, is trained to denoise the data. In modern implementations, this is achieved by training a neural network $\epsilon_{\theta}$ to predict the added noise $\epsilon$, where $\theta$ represents learnable parameters of the neural network, conditioned on the noisy data $x^{t}$, the step $t$, and any contextual information $c$. The objective function is:
@@
In DCM framework, the conditioning context $c$ for a variable $X_{k}$ is its set of parent variables $X_{\mathrm{PA}_{k}}$.
3 Partial Functional Dynamic Backdoor Diffusion-based Causal Model
In SCMs, the presence of unmeasured confounders exhibiting spatial heterogeneity and temporal dependencies renders existing DCMs and BDCMs inadequate for addressing genuine causal inquiries among variables. To surmount this fundamental limitation, we propose a Spatio-Temporal Dynamic Structural Causal Model (ST-DSCM) based on the Backdoor Criterion. We commence by establishing several essential definitions. Due to the complexity of the model, some symbols are mixed up in the article, but the overall readability remains unaffected.
In causal inference, a confounder denotes a variable that causally influences both a treatment variable $X$ and an outcome variable $Y$, thereby inducing a non-causal association between them. Formally, a confounder $Z$ must satisfy: i) $Z$ has a causal effect on $X$; ii) $Z$ has a causal effect on $Y$; iii) $Z$ is not affected by $X$ (i.e., not on the causal pathway between $X$ and $Y$) (Pearl,, 2009).
Observable Confounders refer to measurable variables that satisfy the above conditions and can be measured, which permit adjustment through statistical methods such as stratification, matching or regression (Pearl,, 2009).
Unobservable Confounders denote latent variables that fulfill confounding criteria but resist direct measurement, potentially biasing causal estimates when unaccounted for.
Unobserved explanatory (explained) nodes are unobserved confounder nodes that have no unobserved confounder nodes as its parent (descendant).
3.1 Spatio-Temporal Dynamic Structural Causal Model (ST-DSCM)
Figure 1: Spatio-Temporal Dynamic Structural Causal Model with twelve exogenous and endogenous nodes (where nodes $X_{5}$, $X_{6}$ and $X_{7}$ are unmeasured confounder variables with spatial heterogeneity and temporal dependencies)
Suppose there is a spatio-temporal dataset $\boldsymbol{X}:=\{X_{k}\}_{k\in[K]}$ containing $K$ variables across $n$ regions over $J$ time points. Using $i$ to index regions and $j$ for time points, $X_{k}=(X_{ijk})_{n\times J}$ represent the value of the $k$-th variable $X_{k}$ over the time points and regions. We consider a DAG to characterize the causal relationships among $\{X_{k}\}_{k\in[K]}$.
Consider a DAG $\mathcal{G}$ (e.g. Figure 1) with nodes $[K]$ and a topologically sorted order such that each node $k$ has the $X_{k}$ as the random variable. Let $\mathcal{B}_{k}$ be the set of backdoor nodes [^2] A set of node $\mathcal{B}$ satisfies backdoor criterion (Pearl et al.,, 2016) for tuple $(X,Y)$ in DAG $\mathcal{G}$ if no node in $\mathcal{B}$ is a descendant of $X$ and $\mathcal{B}$ blocks all paths between $X$ (cause) and $Y$ (outcome) which contains an arrow into $X$. of $k$, and $\boldsymbol{X}_{\mathcal{B}_{k}}:=\{X_{l}\}_{l\in\mathcal{B}_{k}}$ represent the variables on $\mathcal{B}_{k}$. Furthermore, let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}\subseteq[K]$ be two distinct sets of nodes with unobserved confounders, where $\mathcal{C}_{1}$ designates a set of unobserved explanatory nodes and $\mathcal{C}_{2}$ denotes a set of unobserved explained nodes. For $h=1,2$, let $\boldsymbol{X}_{\mathcal{C}_{h}}:=\{\boldsymbol{X}_{{\mathcal{C}_{h}}ij}\}_{i\in[n],j\in[J]}:=\{X_{qij}\}_{i\in[n],j\in[J],q\in\mathcal{C}_{h}}$, $\boldsymbol{U}_{\mathcal{C}_{h}}:=\{\boldsymbol{U}_{{\mathcal{C}_{h}}ij}\}_{i\in[n],j\in[J]}:=\{U_{qij}\}_{i\in[n],j\in[J],q\in\mathcal{C}_{h}}$. For example in Figure  1, $\boldsymbol{X}_{\mathcal{C}_{1}}=(X_{5},X_{6})^{\top}$, $\boldsymbol{X}_{\mathcal{C}_{2}}=(X_{7})^{\top}$.
$\boldsymbol{X}_{{\mathcal{C}_{2}}ij}=\boldsymbol{\Gamma}_{i}\boldsymbol{X}_{{\mathcal{C}_{1}}ij}+\boldsymbol{U}_{{\mathcal{C}_{2}}ij},$ (2)
where $\boldsymbol{\Gamma}_{i}$ is a $\dim(\boldsymbol{X}_{{\mathcal{C}_{2}}ij})\times\dim(\boldsymbol{X}_{{\mathcal{C}_{1}}ij})$ structural coefficient matrix, which are permitted to vary across regions.
To capture the temporal dependence, we let $\boldsymbol{X}_{{\mathcal{C}_{1}}i}=(\boldsymbol{X}_{{\mathcal{C}_{1}}i1}^{\top},\cdots,\boldsymbol{X}_{{\mathcal{C}_{1}}iJ}^{\top})^{\top}$ and $\boldsymbol{U}_{{\mathcal{C}_{2}}i}=(\boldsymbol{U}_{{\mathcal{C}_{2}}i1}^{\top},\cdots,\boldsymbol{U}_{{\mathcal{C}_{2}}iJ}^{\top})^{\top}$, then, the covariance matrices $\boldsymbol{\Sigma}_{\boldsymbol{X}_{{\mathcal{C}_{1}}i}}$ and $\boldsymbol{\Sigma}_{\boldsymbol{U}_{{\mathcal{C}_{2}}i}}$ of $\boldsymbol{X}_{{\mathcal{C}_{1}}i}$ and $\boldsymbol{U}_{{\mathcal{C}_{2}}i}$ can be expressed as
$\boldsymbol{\Sigma}_{\boldsymbol{X}_{{\mathcal{C}_{1}}i}}=\boldsymbol{D}_{{\mathcal{C}_{1}}i}\otimes\boldsymbol{T}_{\boldsymbol{X}_{{\mathcal{C}_{1}}}},\ \boldsymbol{\Sigma}_{\boldsymbol{U}_{{\mathcal{C}_{2}}i}}=\boldsymbol{D}_{{\mathcal{C}_{2}}i}\otimes\boldsymbol{T}_{\boldsymbol{U}_{\mathcal{C}_{2}}},$ (3)
where $\boldsymbol{D}_{{\mathcal{C}_{h}}i}(h\in\{1,2\})$ are the $J\times J$ adjacent time covariance matrices, $\boldsymbol{T}_{\boldsymbol{X}_{{\mathcal{C}_{1}}}}$ and $\boldsymbol{T}_{\boldsymbol{U}_{{\mathcal{C}_{2}}}}$ represent between-variable covariances of $\boldsymbol{X}_{{\mathcal{C}_{1}}ij}$ and $\boldsymbol{U}_{{\mathcal{C}_{2}}ij}$, respectively.
By incooperating the spatio-temporal dynamic structure among unmeasured confounders into the traditional SCM framework, we develop a novel Spatio-Temporal Dynamic Structural Causal Model (ST-DSCM) to describe the relationship between an observed node $k$ and its causal backdoor nodes $\boldsymbol{X}_{\mathcal{B}}$ with unmeasured confounders $\{\boldsymbol{X}_{\mathcal{C}_{h}}\}_{h\in\{1,2\}}$ and $\boldsymbol{U}_{\mathcal{C}_{2}}$ exhibit spatio-temporal dynamic structural relationships as constructed in the above.
However, real-world applications often present significant challenges in elucidating causal relationships among variables observed under heterogeneity. To address these complexities, we extend the ST-DSCM by incorporating functional random variables $\boldsymbol{X}_{k}(t)$, leading to a Partial Functional Spatio-Temporal Dynamic Structural Causal Model (PFST-DSCM), as described in Figure 2.
3.2 PFST-DSCM: Partial Functional Spatio-Temporal Dynamic Structural Causal Model
We adopt a basis expansion framework to achieve dimensionality reduction in the functional space (Zoh et al.,, 2024) via a set of orthogonal basis functions $\{\mathbf{b}_{ij1},\cdots,\mathbf{b}_{ijK_{n}}\}\in\mathbb{R}^{T\times K_{n}}$. Let
$X_{ijm}=\int\mathbf{b}_{ijm}(t)X_{ij}(t)dt,$ (4)
for $i\in[n],j\in[J],m\in[K_{n}]$, which are used as nodes within the ST-DSCM (see Figure 2).
Figure 2: Partial Functional Spatio-Temporal Dynamic Structural Causal Model with 33 exogenous and endogenous nodes (where nodes $X_{28}$, $X_{29}$ and $X_{30}$ are unmeasured confounders with spatial heterogeneity and temporal dependencies, $Y_{1}(t),Y_{2}(t),Y_{3}(t)$ are functional nodes, and $X_{4},\cdots,X_{21}$ are the corresponding base expansion nodes)
We assume that the unobserved random variables are jointly independent (Markovian SCM), and the Spatio-Temporal Dynamic Directed Acyclic Graph (ST-DDAG) $\mathcal{G}$ is the graph induced by ST-DSCM $\mathcal{M}$. Every ST-DSCM $\mathcal{M}$ entails a unique joint observational distribution satisfying the causal Markov assumption: $p(\mathbf{X})=\prod_{k=1}^{K}p(X_{k}|X_{\mathcal{B}_{k}})$.
3.3 PFD-BDCM
We introduce the PFD-BDCM for modeling PFST-DSCM and addressing causal queries. The PFD-BDCM approach leverages on an encoder-decoder architecture to construct a PFST-DSCM, which facilitates elucidation of causal relationships among variables observed at varying resolutions in the presence of unmeasured confounders that exhibit spatial heterogeneity and temporal dependencies.
When unmeasured confounders exist with spatial heterogeneity and temporal dependencies, we model a data generating process for observed variables $\{\mathbf{X}_{k}:X_{ijk}=f_{ij}(\mathbf{X}_{\mathcal{B}_{k}},\mathbf{U}_{k})\}_{k\in[K]}$.
Our framework has an encoding function $g$ and a decoding function $h$. The encoder $g$ maps the observed data $(\mathbf{X}_{k},\mathbf{X}_{\mathcal{B}_{k}})$ to a latent variable $\mathbf{Z}_{k}$, defined as $\mathbf{Z}_{k}:=g(\mathbf{X}_{k},\mathbf{X}_{\mathcal{B}_{k}})$, aiming to encapsulate information pertaining to $\mathbf{U}_{k}$. The decoder $h$ reconstructs $\mathbf{X}_{k}$ from $(\mathbf{Z}_{k},\mathbf{X}_{\mathcal{B}_{k}})$, yielding $\hat{\mathbf{X}}_{k}=h(\mathbf{Z}_{k},\mathbf{X}_{\mathcal{B}_{k}})$. Under ideal conditions, perfect reconstruction ($\hat{\mathbf{X}}_{k}=\mathbf{X}_{k}$) implies that $h$ effectively approximates the true structural function $f$.
We first elaborate on the architecture and training procedure of PFD-BDCM, and then explain how to employ it for answering causal queries. We start with some notations.
  • Let $Z_{k}^{t}$ be the $k$-th exogenous node value at diffusion step $t$ of the forward implicit diffusion process and let $Z_{k}\coloneqq Z_{k}^{T}$, where $T$ represent the total diffusion steps.
  • Let $\hat{X}_{k}^{t}$ be the $k$-th endogenous node value at diffusion step $t$ of the reverse implicit diffusion process and let $\hat{X}_{k}\coloneqq\hat{X}_{k}^{0}$.
@@ 3.3 PFD-BDCM
Causal inference, particularly counterfactual reasoning, needs a deterministic mapping from observational data to latent representations (encoding) and subsequent reconstruction (decoding). While standard Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al.,, 2020) employ stochastic reverse processes, Denoising Diffusion Implicit Models (DDIMs) (Song et al.,, 2021) use a non-Markovian formulation that renders the reverse transition deterministic—a property of fundamental importance. Within the PFD-BDCM framework, the encoding and decoding procedures extend the DDIM paradigm by integrating backdoor adjustment sets as supplementary covariates, and the diffusion model for node $k$ is denoted by $\epsilon_{\theta}^{k}(X_{k},\mathbf{X}_{\mathcal{B}_{k}},t)$.
@@
$Z^{t+1}_{k}$= $\sqrt{\alpha_{t+1}/\alpha_{t}}Z^{t}_{k}+\epsilon_{\theta}^{k}(Z^{t}_{k},\mathbf{X}_{\mathcal{B}_{k}},t)$ (5)
$\times(\sqrt{1-\alpha_{t+1}}-\sqrt{\alpha_{t+1}(1-\alpha_{t})/\alpha_{t}}),$
@@
$\hat{X}^{t-1}_{k}$= $\sqrt{\alpha_{t-1}/\alpha_{t}}\hat{X}^{t}_{k}-\epsilon_{\theta}^{k}(\hat{X}_{k}^{t},\mathbf{X}_{\mathcal{B}_{k}},t)$ (6)
$\times\left(\sqrt{\alpha_{t-1}(1-\alpha_{t})/\alpha_{t}}-\sqrt{1-\alpha_{t-1}}\right),$
for $t=T,\cdots,1$, initialized with $\hat{X}^{T}_{k}:=Z_{k}$. The encoding and decoding functions for node $k$ are denoted as $\operatorname{Enc}_{k}(X_{k},\mathbf{X}_{\mathcal{B}_{k}})$ (Eq. (5)) and $\operatorname{Dec}_{k}(Z_{k},\mathbf{X}_{\mathcal{B}_{k}})$ (Eq. (6)) respectively, whose pseudocodes are provided in Appendix A.1.
@@
The comprehensive training methodology (Algorithm 1) incorporates backdoor adjustment sets as covariates while training distinct diffusion models per node. Crucially, generative models for endogenous nodes exhibit mutual independence during training, thereby enabling parallelized optimization. This parallelism is feasible since each diffusion model necessitates only its target node’s values and corresponding backdoor adjustment set values. The final PFD-BDCM architecture integrates these $K$ trained diffusion models $\{\epsilon_{\theta}^{k}\}_{k\in[K]}$.
@@
Generating samples for observational/interventional queries. To generate samples approximating the interventional distribution $p(\mathbf{X}|\mathrm{do}(X_{\mathcal{L}}:=\boldsymbol{\gamma}))$ using a trained PFD-BDCM model, we implement the following procedure: i) For intervened nodes $l\in\mathcal{L}$, we set $\hat{X}_{l}:=\gamma_{l}$ deterministically; ii) For root nodes $k$, sample $\hat{X}_{k}$ from empirical training distributions; iii) For non-intervened nodes $k\notin\mathcal{L}$, sample latent vectors $Z_{k}\sim\mathcal{N}(\mathbf{0},I_{d_{k}})$, where $d_{k}=\dim(X_{k})$, and subsequently compute $\hat{X}_{k}:=\mathrm{Dec}_{k}(Z_{k},\hat{\mathbf{X}}_{\mathcal{B}_{k}})$ utilizing inductively generated backdoor variable values $\hat{\mathbf{X}}_{\mathcal{B}_{k}}$. Generated values propagate to child nodes as backdoor inputs. Observational sampling ($p(\mathbf{X})$) corresponds to $\mathcal{L}=\emptyset$, with pseudocode formalized in Algorithm 2.
Algorithm 1 PFD-BDCMs Training
Input: target distribution $\mathcal{Q}$, scale factors $\{\alpha_{t}\}_{t=1}^{T}$, ST-DDAG $\mathcal{G}$ node $k$ is represented by $X_{k}$ and intervened node $l$ with intervened value $\gamma_{l}$}
1:while not converged do
2: Sample $X_{k}^{0}\sim\mathcal{Q}$
3:for $k=1,\cdots,K$ do
4:  $t\sim\mathrm{Unif}[\{1,\cdots,T\}]$, $\varepsilon\sim\mathcal{N}(0,I)$
5:  Update the parameter of the node $k$ ’s diffusion model $\epsilon_{\theta}^{k}$ by minimizing the Eq. (1) depending on the nodes.
6:  if $X_{l}\in\mathbf{X}_{\mathcal{B}_{k}}$ then
7:   $\left\|\epsilon-\epsilon_{\theta}^{k}\left(\sqrt{\alpha_{t}}X_{k}^{0}+\sqrt{1-\alpha_{t}}\epsilon,\mathbf{X}_{\mathcal{B}_{k}}^{0},X_{l},t\right)\right\|_{2}^{2}$
8:  else
9:   $\left\|\epsilon-\epsilon_{\theta}^{k}\left(\sqrt{\alpha_{t}}X_{k}^{0}+\sqrt{1-\alpha_{t}}\epsilon,\mathbf{X}_{\mathcal{B}_{k}}^{0},t\right)\right\|_{2}^{2}$
10:  end if
11:end for
12:end while
Algorithm 2 PFD-BDCMs Observational/Interventional Sampling
Input: Intervened node $l$ with value $\gamma_{l}$ ($\mathcal{L}=\emptyset$ for observational sampling).
1:for $k=1,\cdots,K\{\text{in topological order}\}$  do
2:if  $k=l$  then
3:  $\hat{X}_{k}\leftarrow\gamma_{l}$
4:else if $k$ is a root node then
5:  $\hat{X}_{k}\sim E_{k}$
6:else if $X_{l}\in\mathbf{X}_{\mathcal{B}_{k}}$ then
7:  $\hat{X}_{k}\leftarrow\operatorname{Dec}_{k}(Z_{k},\hat{\mathbf{X}}_{\mathcal{B}_{k}},X_{l})$
8:else
9:  $\hat{X}_{k}\leftarrow\operatorname{Dec}_{k}(Z_{k},\hat{\mathbf{X}}_{\mathcal{B}_{k}})$
10:end if
11:end for
[…]
@@
Counterfactual Queries. To compute counterfactual estimates $\hat{\mathbf{x}}^{\mathrm{CF}}$ within the PFD-BDCM framework, given factual observation $\mathbf{x}^{\mathrm{F}}:=(x_{1}^{\mathrm{F}},\cdots,x_{K}^{\mathrm{F}})$ and intervention set $\mathcal{L}$ with values $\boldsymbol{\gamma}$, we implement the following systematic procedure: i) For intervened nodes $l\in\mathcal{L}$, assign $\hat{x}_{l}^{\mathrm{CF}}:=\gamma_{l}$ deterministically; ii) For non-intervened descendant nodes $k$, using inductively generated backdoor estimates $\hat{\mathbf{x}}_{\mathcal{B}_{k}}^{\mathrm{CF}}$, we compute $\hat{x}_{k}^{\mathrm{CF}}:=\mathrm{Dec}_{k}(\mathsf{Enc}_{k}(x_{k}^{\mathrm{F}},\mathbf{x}_{\mathcal{B}_{k}}^{\mathrm{F}}),\hat{\mathbf{x}}_{\mathcal{B}_{k}}^{\mathrm{CF}})$, where factual noise is implicitly encoded. The complete formalization appears in Appendix A.2.
4 Bounding Counterfactual Error

### v2

@@ 2.2 Partially Functional Dynamic Backdoor Diffusion-based Causal Model
Our approach leverages the generative power of diffusion models to learn the complex functional relationships inherent in a Structural Causal Model. The core idea is to represent the structural equation for each endogenous variable $X_{k}$ with a dedicated conditional diffusion model. This model learns the distribution $p(X_{k}|X_{\mathrm{PA}_{k}},U_{k})$ . Diffusion models (Sohl-Dickstein et al.,, 2015; Ho et al.,, 2020) approximate a target data distribution $q(x^{0})$ via a two-stage process. First, a fixed forward process gradually injects Gaussian noise into the data $x^{0}$ over $T$ steps. The distribution of the noisy data $x^{t}$ at step $t$ is a Gaussian: $q(x^{t}|x^{0})=\varphi(x^{t};\sqrt{\alpha_{t}}x^{0},(1-\alpha_{t})I)$, where ${\varphi}(x;\mu,\Sigma)$ denote the Gaussian density with mean $\mu$ and covariance $\Sigma$. Here, $\alpha_{t}:=\prod_{s=1}^{t}(1-\beta_{s})$ where $\beta_{s}$ is a predefined noise schedule at each time step $s$. As $t\to T$, $x^{T}$ converges to a standard Gaussian distribution. Second, a learnable reverse process, parameterized by $\theta$, is trained to denoise the data. This is achieved by training a neural network $\epsilon_{\theta}$ to predict the added noise $\epsilon$ using the available observational data. For learning the distribution $p(X_{k}|X_{\mathrm{PA}_{k}},U_{k})$, the network $\epsilon_{\theta}$ is trained on samples where $X_{k}$ serves as the target variable $x^{0}$ and $X_{\mathrm{PA}_{k}}$ provides the conditioning context $c$. Here $\theta$ represents learnable parameters of the neural network, conditioned on the noisy data $x^{t}$, the step $t$, and the parental variables contextual information $c$. The objective function is
@@
In our framework, the conditioning context $c$ for a variable $X_{k}$ is its set of parent variables $X_{\mathrm{PA}_{k}}$ (or $X_{\mathcal{B}_{k}}$).
@@
Although diffusion models excel at data generation, causal inference, especially counterfactual reasoning, requires a deterministic mapping between observations and latent codes. Denoising Diffusion Implicit Models (DDIMs) (Song et al.,, 2021) provide such a deterministic non-Markovian reverse process, a property essential for identifiability. In PFD-BDCM, we extend DDIM by incorporating backdoor adjustment sets as additional covariates. The resulting diffusion model for node $k$ is denoted $\epsilon_{\theta}^{k}(X_{k},\mathbf{X}_{\mathcal{B}_{k}},t)$.
@@
$Z^{t+1}_{k}$= $\sqrt{\alpha_{t+1}/\alpha_{t}}Z^{t}_{k}+\epsilon_{\theta}^{k}(Z^{t}_{k},\mathbf{X}_{\mathcal{B}_{k}},t)(\sqrt{1-\alpha_{t+1}}-\sqrt{\alpha_{t+1}(1-\alpha_{t})/\alpha_{t}}),$ (6)
@@
$\hat{X}^{t-1}_{k}$= $\sqrt{\alpha_{t-1}/\alpha_{t}}\hat{X}^{t}_{k}-\epsilon_{\theta}^{k}(\hat{X}_{k}^{t},\mathbf{X}_{\mathcal{B}_{k}},t)\left(\sqrt{\alpha_{t-1}(1-\alpha_{t})/\alpha_{t}}-\sqrt{1-\alpha_{t-1}}\right),$ (7)
for $t=T,\cdots,1$, initialized with $\hat{X}^{T}_{k}:=Z_{k}$. The encoding and decoding functions for node $k$ are denoted as $\mathrm{Enc}_{k}(X_{k},\mathbf{X}_{\mathcal{B}_{k}})$ (Eq. (6)) and $\mathrm{Dec}_{k}(Z_{k},\mathbf{X}_{\mathcal{B}_{k}})$ (Eq. (7)) respectively, and the pseudocodes are provided in Appendix B.1.
@@
The comprehensive training methodology (Appendix B.2 Algorithm 3 ) incorporates backdoor adjustment sets as covariates while training distinct diffusion models per node. Crucially, generative models for endogenous nodes exhibit mutual independence during training, thereby enabling parallelized optimization. This parallelism is feasible since each diffusion model necessitates only its target node’s values and corresponding backdoor adjustment set values. The final PFD-BDCM architecture integrates these $K$ trained diffusion models $\{\epsilon_{\theta}^{k}\}_{k\in[K]}$.
@@
To generate samples approximating the interventional distribution $p(\mathbf{X}|\mathrm{do}(X_{\mathcal{L}}:=\boldsymbol{\gamma}))$ using a trained PFD-BDCM model, we implement the following procedure: i) For intervened nodes $l\in\mathcal{L}$, we set $\hat{X}_{l}:=\gamma_{l}$ deterministically; ii) For root nodes $k$, sample $\hat{X}_{k}$ from empirical training distributions; iii) For non-intervened nodes $k\notin\mathcal{L}$, sample latent vectors $Z_{k}\sim\mathcal{N}(\mathbf{0},I_{d_{k}})$, where $d_{k}=\dim(X_{k})$, and subsequently compute $\hat{X}_{k}:=\mathrm{Dec}_{k}(Z_{k},\hat{\mathbf{X}}_{\mathcal{B}_{k}})$ utilizing inductively generated backdoor variable values $\hat{\mathbf{X}}_{\mathcal{B}_{k}}$. Generated values propagate to child nodes as backdoor inputs. Observational sampling ($p(\mathbf{X})$) corresponds to $\mathcal{L}=\emptyset$, with pseudocode formalized in Appendix B.3 (Algorithm 4 ).
Counterfactual Queries.
@@
To compute counterfactual estimates $\hat{\mathbf{x}}^{\mathrm{CF}}$ within the PFD-BDCM framework, given factual observation $\mathbf{x}^{\mathrm{F}}:=(x_{1}^{\mathrm{F}},\cdots,x_{K}^{\mathrm{F}})$ and intervention set $\mathcal{L}$ with values $\boldsymbol{\gamma}$, we implement the following systematic procedure: i) For intervened nodes $l\in\mathcal{L}$, assign $\hat{x}_{l}^{\mathrm{CF}}:=\gamma_{l}$ deterministically; ii) For non-intervened descendant nodes $k$, using inductively generated backdoor estimates $\hat{\mathbf{x}}_{\mathcal{B}_{k}}^{\mathrm{CF}}$, we compute $\hat{x}_{k}^{\mathrm{CF}}:=\mathrm{Dec}_{k}(\mathrm{Enc}_{k}(x_{k}^{\mathrm{F}},\mathbf{x}_{\mathcal{B}_{k}}^{\mathrm{F}}),\hat{\mathbf{x}}_{\mathcal{B}_{k}}^{\mathrm{CF}})$, where factual noise is implicitly encoded. The complete formalization appears in Appendix B.3 (Algorithm 5 ).
3 Counterfactual Error Bounds